OpenStudy (anonymous):

find dr/dt of sin(rt) = (1/5)

OpenStudy (anonymous):

using implicit differentiation

myininaya (myininaya):

\[(rt)'\cos(rt)=0=>(r't+r)\cos(rt)=0\]

myininaya (myininaya):

\[r't \cos(rt)+r \cos(rt)=0\]

myininaya (myininaya):

\[r' t \cos(rt)=-r \cos(rt)\] \[r' =\frac{-r \cos(r t)}{t \cos(r t)}\] \[r'=\frac{-r}{t}, \cos(rt) \neq 0\]

OpenStudy (anonymous):

THANK YOU!

myininaya (myininaya):

np :)